40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 23     Entry time: Mon Oct 29 09:16:31 2007
Author: steve 
Type: Routine 
Category: VAC 
Subject: the rga is back 
We had no filament current since last power glitch of Oct. 8, 2007
First I thought that the filament was lost, but it was only bad contact.
The rga head pins were oxidized. Rga was turned back on last Friday.
It's temp is 55.3C normal
Attachment 1: rgaisback.jpg  273 kB  | Hide | Hide all
rgaisback.jpg
ELOG V3.1.3-